# A Convex-Inspired Neural Construction for Structured and Generalizable Nonlinear Model Reduction

### v1

@@ 5.2. Optimization with Small Amount of Cubature Points
Figure 10. We train each method on only five keyframes from a collapsing camel sequence and test their ability to interpolate and generalize under sparse supervision. Due to the limited data, the linear basis is constrained to five components, while nonlinear models use eight latent coordinates. Under latent-space perturbations, the linear model exhibits locally degenerate behavior, and the vanilla autoencoder produces unstable, implausible deformations. In contrast, our convex-inspired formulation maintains smooth and coherent reconstructions, demonstrating improved robustness and physically meaningful interpolation even with extremely limited training data.
@@
5.3. Generalization on Sparsely Sampled Training Data
Our method can also be applied to interpolate deformation data after training on keyframes, demonstrating improved robustness when the training set is small. As shown in Fig. 10, we evaluate generalization under a supervised setting with extremely sparse training data, consisting of only five key frames from the collapse sequence. Such sparsity constrains the linear PCA basis to five components, while nonlinear models employ eight latent coordinates to capture higher-order deformations. Despite this limited supervision, our architecture achieves comparable reconstruction error on the training frames and smaller residuals on unseen configurations. When perturbing the latent code $\bm{q}$ with large Gaussian noise $\epsilon\sim\mathcal{N}(0,15^{2}I)$, the vanilla MLP mapping produces unstable and physically implausible states, whereas the linear mapping exhibits locally degenerate regions due to its restricted spectral capacity. In contrast, our architecture maintains smooth and coherent reconstructions across a wide range of latent perturbations. This robustness arises from the convexity-inspired inductive bias, which regularizes the latent-to-physical mapping, mitigates overfitting under sparse supervision, and promotes physically meaningful interpolation across the under-sampled deformation manifold.
5.4. Collision Handling
@@
5.5. Real-time Simulations
@@ 6. Discussions and Future Work
We demonstrate the main advantage of our method through its stronger generalization on unseen data—both when exploring the latent reduced space and during deformation interpolation between training snapshots. Within regions well covered by training data, our formulation exhibits convergence behavior comparable to that of a vanilla autoencoder. As shown in Figure 13, we train on two different shapes and minimize their elastic energy until returning to rest, starting from different coordinates in the reduced space. The average normalized elastic energy and its variance are used to quantify convergence.
Thanks to the compactness of the nonlinear representation, our method achieves comparable deformation accuracy with fewer basis functions, resulting in a smaller system and faster solves than linear reduction. When compared to other nonlinear models, our formulation converges faster in some cases and slightly slower in others, suggesting that convergence may depend on the underlying shape family or energy landscape.

### v2

@@ 5.4. Optimization with Small Amount of Cubature Points
5.5. Real-time Simulations
@@
5.6. Discretization-agnostic Extension
We further demonstrate that our regularization method extends beyond mesh-based model reduction to continuous nonlinear settings. Specifically, we apply it to the CROM framework (Chen et al., 2023b), a discretization-agnostic reduced model based on implicit neural representations. As shown in Figure 15, we train a model to simulate deformations of an elephant head under a single poking force. We then evaluate generalization under scaled and inverted forces. While CROM performs well within the training regime, it struggles to extrapolate: it fails to scale with increased force magnitude and cannot invert the deformation under force reversal. In contrast, our method generalizes to both scenarios, producing physically plausible responses under unseen loading conditions.
To demonstrate that our method preserves the discretization-agnostic property of prior continuous model reduction approaches, we evaluate a single trained model on meshes of three different resolutions: 2.5K, 50K, and 250K vertices. Despite the varying discretization, the model produces qualitatively similar deformation behavior across all resolutions, indicating robust generalization to unseen mesh discretizations.
Lastly, we demonstrate that the continuous formulation of our method can reconstruct plausible deformations from sparse, real-world observations. We train a deformation model on a rubber toy using only 2D keypoint annotations extracted from a short video sequence (Fig. 17). Despite the limited supervision, the model accurately reconstructs full-object deformations and generalizes to new frames involving unseen forces. This example highlights the potential of our formulation and regularization in real-world settings, supporting data-efficient deployment without requiring dense input.
[…]
Figure 16. Discretization-agnostic generalization. A single model, trained on a continuous representation, is evaluated on three meshes with increasing resolution (#V: 2.5K, 50K, 250K). The resulting deformations remain consistent across discretizations, demonstrating that our method inherits the resolution-independence of prior continuous model reduction techniques.
Figure 17. Real-world deformation reconstruction using our continuous formulation and sparse 2D keypoint supervision. (1) A video of a rubber toy undergoing deformation is recorded. (2) A sparse set of 2D keypoints is annotated for a small number of frames. (3) The model is trained to reconstruct full-object deformations from these annotations. (4) Our method generalizes to reconstruct plausible deformations under unseen forces, despite being trained only on limited keypoint data.
@@ 6. Discussions and Future Work
We demonstrate the main advantage of our method through its improved generalization to unseen loading conditions, including both increased force magnitudes and reversed directions. This capability stems from our structured decoder, which integrates convexity and odd symmetry to better capture the physical behavior of deformable systems under out-of-distribution forces.
Due to the compactness of the nonlinear representation, our method achieves comparable deformation accuracy using significantly fewer basis functions, resulting in smaller reduced systems and faster solves compared to traditional linear reduction.
Within regions well covered by training data, our formulation exhibits convergence behavior similar to that of a vanilla autoencoder—converging faster in some cases and slightly slower in others. As shown in Figure 18, we evaluate convergence by minimizing elastic energy from various starting points in the reduced space and measure progress via normalized energy and its variance over iterations.